# On the connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity

### v1

@@ II.1.1 Partial ordering and Möbius Inversion
The incidence algebra is a unital algebra with the identity element defined as $\mathrm{I}(x,y)=\delta_{xy}$, where $\delta_{xy}$ is the usual Kronecker delta. With the identity being present in the incidence algebra, the inverse of $f\in R[S\times S]$ can be defined if there exist $f^{-1}\in R[S\times S]$ such $f*f^{-1}(x,y)=f^{-1}*f(x,y)=\mathrm{I}(x,y)$ for all $x,y\in S$ with $x\leq y$. Equivalently
@@
The above equation is used to find $f^{-1}(x,y)$ inductively. Specifically, for all $x\in S$, we have $f^{-1}(x,x)=\frac{1}{f(x,x)}$ from the above equation. Then this data is used to find $f^{-1}(x,z)$ for $z\in S$ such that there doesn’t exist $z_{0}\in S$ except $x$ and $z$ with $x\leq z_{0}\leq z$. Continuing in this way, it is possible to find $f^{-1}(x,y)$.
Definition.
(Zeta function and Möbius function) The zeta function is an element $\zeta$ in $R[S\times S]$ such that $\zeta(x,y)=1$ for all $x\leq y$ and 0 otherwise. The Möbius function denoted by $\mu$ is the inverse of the zeta function.
It is clear that $\mu(x,x)=\frac{1}{\zeta(x,x)}=1$ and when $x<y$, from Eq. (2), we have
@@
The above equation can be used inductively to determine the Möbius function.
Theorem B.
(Möbius Inversion) Let $S$ be a partially ordered set with the ordering denoted by $\leq$ and let $R$ be a field. Let $f:S\to R$ and $h:S\to R$ be two functions satisfying for all $x\in S$
@@
Then $h$ can be expressed in terms of $f$ as follows:
@@
II.1.2 Steinberg’s Isomorphism
An example of an inverse semigroup is the symmetric inverse semigroup. Denoted by $\mathcal{J}_{X}$, it is the semigroup consists of all partial bijections $\tau:Y\to Z$ on a non-empty set $X$, where $X,Y\subseteq X$. The binary operation is the composition of partial transformations. It can be verified that the $\tau:Y\to Z$ and $\tau^{-1}:Z\to Y$ satisfies all the defining properties of an abstract inverse semigroup defined in Definition (A). Now, the Preston-Vagner representation theorem [3] says that any inverse semigroup $S$ can be thought of as an inverse subsemigroup of $\mathcal{J}_{X}$.
Preston-Vagner representation: Let $S$ be an inverse semigroup and $\mathcal{J}_{S}$ be the symmetric inverse semigroup of all partial bijection of $S$. Then $S$ is isomorphic to an inverse subsemigroup of $\mathcal{J}_{S}$.
For $\tau\in\mathcal{J}_{S}$, we denote by $\mathrm{dom}(\tau)$ the domain of $\tau$ and by $\mathrm{ran}(\tau)$ the range of $\tau$. Denoting by $\mathrm{id}_{U}$ the identity map on the subset $U\subseteq S$, we have $\tau\tau^{-1}=\mathrm{id}_{\mathrm{ran}(\tau)}$ and $\tau^{-1}\tau=\mathrm{id}_{\mathrm{dom}(\tau)}$. As $\mathrm{id}_{U}$ can be identified with the subset $U\subseteq S$ itself, it a customary to define for an abstract inverse semigroup $S$ and $s\in S$
$\mathrm{dom}(s):=s^{-1}s,$ (6)
$\mathrm{ran}(s):=ss^{-1}$ (7)
and think of $s$ as an isomorphism $s:s^{-1}s\to ss^{-1}$. Recalling that $s^{-1}s$ and $ss^{-1}$ are idempotents, the notion of two idempotents being isomorphic is defined as follows:
Definition C.
(Isomorphic idempotents)[4] Let $S$ be an inverse semigroup. The idempotents $e,e^{\prime}\in S$ are said to be isomorphic if there exists $s\in S$ such that $e=s^{-1}s$ and $e^{\prime}=ss^{-1}$.
@@ II.1.2 Steinberg’s Isomorphism
Definition D.
Let $S$ be an inverse semigroup and $s,t\in S$. Then $s\leq t$ if there exist an idempotent $e\in S$ such that $s=et$.
For two arbitrary elements $s$ and $t$ belonging to an inverse semigroup $S$, the following are equivalent.
  1. 1.
    $s\leq t$.
  2. 2.
    There exists and idempotent $e^{\prime}$ such that $s=te^{\prime}$.
  3. 3.
    $s=ss^{-1}t$.
  4. 4.
    $s=ts^{-1}s$.
Definition E.
[4] Let $S$ be a finite inverse semigroup. For every non-zero element $s\in S$, let us define the following:
$\lfloor s\rfloor:=\sum_{t\in S,t\leq s}\mu(t,s)t\in\mathbb{C}_{0}[S],$ (8)
where $\mu(t,s)$ is the Möbius function defined on the partial order of $S$ as given in Eq. (3).
It simply follows that for the zero element $z\in S$, $\lfloor z\rfloor=0$. It is proved in Ref. [4], that $\{\lfloor s\rfloor\ :s\neq z,s\in S\}$ is a basis for the contracted algebra $\mathbb{C}_{0}[S]$. The basis $\{\lfloor s\rfloor\}_{s\in S\setminus z}$ is known as the groupoid basis. Using the Möbius inversion formula, we can recover the natural basis $\{s\}_{s\in S\setminus z}$ as :
$s=\sum_{t\in S,t\leq s}\lfloor t\rfloor.$ (9)
The basis elements in $\{\lfloor s\rfloor\}_{s\in S\setminus z}$ multiplies as [4]:
@@
Let $S$ be a finite inverse semigroup. Then, $s,t\in S$, are said to be $\mathcal{D}$-relatable and denote by $s\mathcal{D}t$ if there exist an element $x\in S$ such that $x^{-1}x=ss^{-1}$ and $xx^{-1}=tt^{-1}$
@@
The above defined relation is called Green’s $\mathcal{D}$-relation, and it can be shown that it is an equivalence relation on $S$. The equivalence classes with respect to this relation are known as the $\mathcal{D}$-classes.
@@
Let $\{D_{k}\}_{k=1}^{n}$ be the $\mathcal{D}$-classes of $S$. Let us define $\mathbb{C}D_{k}:=\mathrm{Span}\{\lfloor s\rfloor\}$ where $s\in D_{k}$. It should be noted that there always exists an idempotent inside each $\mathcal{D}$-classes $D_{k}$, because if $t\in D_{k}$ then $t^{-1}t\in D_{k}$. It follows that if $e$ and $f$ are two idempotents in $D_{k}$ then they are isomorphic, i.e., there exists $x\in S$ such that $x^{-1}x=e$ and $xx^{-1}=f$.
Proposition F.
$\mathbb{C}D_{k}$ is an algebra.
@@
Proof: From the definition, $\mathbb{C}D_{k}$ is a vector space. Let $s,t\in D_{k}$ be two arbitrary elements. So, the multiplication of $\lfloor s\rfloor$ and $\lfloor t\rfloor$ is given by Eq. (10). Now, it is sufficient to show that $st\in D_{k}$. As $s,t\in D_{k}$, there exist $x\in S$, such that $x^{-1}x=ss^{-1}$ and $xx^{-1}=tt^{-1}$. Now
$(st)(st)^{-1}$ $=stt^{-1}s^{-1}$
$=ss^{-1}ss^{-1}$
$=ss^{-1}$
$=x^{-1}x,$
@@
where in the 2nd equality, we have used the condition in Eq. (10). It is clear from the above equation that $(st)\mathcal{D}t$ and hence $st\in D_{k}$. This completes the proof. $\Box$
@@
A subset $G\subseteq S$ is called a subgroup of a semigroup $S$ if $G$ is a group with respect to the operations in $S$. A subgroup $G$ of $S$ is called a maximal subgroup if it is not properly contained in some subgroup of $S$ except itself.
Proposition.
@@
[4] Let $S$ be a finite inverse semigroup and $e\in S$ be an idempotent. Then $e$ is the identity element of a unique maximal subgroup $G_{e}\subseteq S$ given by
$G_{e}=\{s\in S\quad|\quad ss^{-1}=s^{-1}s=e\},$
Proposition.
@@
Let $e$ and $f$ be two isomorphic idempotents of a finite inverse semigroup $S$ and let $G_{e}$ and $G_{f}$ represents their maximal subgroups. Then $G_{e}$ is isomorphic to $G_{f}$.
@@
From Proposition (F) and the multiplication rule given in Eq. (10), it follows that $\mathbb{C}_{0}[S]=\bigoplus_{k=1}^{n}\mathbb{C}D_{k}$. For each $D_{k}$, let us fix an idempotent $e_{k}$. Let $G_{e_{k}}$ denotes the maximal subgroup of $S$ at $e_{k}$. For each idempotent $e\in D_{k}$, let us fix an element $p_{e}\in S$ such that $\mathrm{dom}(p_{e})=e_{k}$ and $\mathrm{ran}(p_{e})=e$ with $p_{e_{k}}=e_{k}$. The following theorem states that $\mathbb{C}D_{k}$ is isomorphic to $M_{r_{k}}(\mathbb{C}G_{e_{k}})$ where $M_{r_{k}}(\mathbb{C}G_{e_{k}})$ is the algebra of $r_{k}\times r_{k}$ matrices over the group algebra $\mathbb{C}G_{e_{k}}$ with $r_{k}$ being the number of idempotents in $D_{k}$.
Theorem G.
[4] Let $S$ be a finite inverse semigroup and $\{D_{k}\}_{k=1}^{N}$ represents its $\mathcal{D}$-classes. Let $e_{k}$ be an arbitrary but fixed idempotent in $D_{k}$ and $G_{e_{k}}$ represents the maximal subgroup of at $e_{k}$. Then $\mathbb{C}_{0}[S]\cong\bigoplus_{k=1}^{N}M_{r_{k}}(\mathbb{C}G_{k})$. Specifically, The algebra $\mathbb{C}D_{k}$ is isomorphic to the algebra $M_{r_{k}}(\mathbb{C}G_{k})$, where the isomorphism is given by
$\varphi(\lfloor s\rfloor)=p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}E_{\mathrm{ran}(s),\mathrm{dom}(s)},$ (11)
@@
$\varphi^{-1}(sE_{t,f})=\lfloor p_{t}sp^{-1}_{f}\rfloor.$ (12)
Proposition H.
Let $s\in D_{k}$. Then $p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\in G_{e_{k}}$.
Proof: We begin with the following,
@@
In a similar way, we can also prove that $\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)^{-1}=e_{k}$. This completes the proof. $\Box$
Proposition I.
For all $u\in G_{e_{k}}$, we have $\mathrm{ran}(p_{a}\,u\,p^{-1}_{b})=a$ and $\mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b$.
Proof:
${ran}(p_{a}\,u\,p^{-1}_{b})$ $=p_{a}\,u\,p^{-1}_{b}\,p_{b}\,u^{-1}\,p^{-1}_{a}$
@@
In a similar way, we can prove $\mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b$. This completes the proof. $\Box$
Steinberg’s isomorphism provides a way to construct a complete set of inequivalent irreducible representations of the contracted algebra $\mathbb{C}_{0}[S]$ by using complete sets of inequivalent irreducible representations of the maximal subgroups of the finite inverse semigroup $S$.
Theorem J.
[4] Let $S$ be a finite inverse semigroup and $\{D_{k}\}_{k=1}^{N}$ represents its $\mathcal{D}$-classes. Let $e_{k}$ be an arbitrary but fixed idempotent in $D_{k}$ and $G_{e_{k}}$ represents the maximal subgroup of at $e_{k}$. Let $\mathrm{Irr}(G_{k})$ denotes a complete set of inequivalent irreducible representations of the maximal subgroup $G_{e_{k}}$. For $\rho_{k}\in\mathrm{Irr}(G_{k})$, define a representation $\overline{\rho}_{k}$ on $M_{r_{k}}(\mathbb{C}G_{k})$ by $\overline{\rho}_{k}(gE_{a,b}):=E_{a,b}\otimes\rho_{k}(g)$ with the linear extension to the whole algebra. Now define $\overline{\rho}_{k}$ on the algebra $\bigoplus_{k=1}^{N}M_{r_{k}}(\mathbb{C}G_{k})$ by defining its action on the summands other than $k$ as zero. Then the set of $\overline{\rho}_{k}$ defined in this way forms a complete set of inequivalent irreducible representation of $\mathbb{C}_{0}[S]$.
II.2 CJKS Isomorphism and completely positive maps
Let $\mathcal{H}_{1}=\mathbb{C}^{n}$ and $\mathcal{H}_{2}=\mathbb{C}^{m}$ be two finite-dimensional Hilbert spaces and $B(\mathcal{H}_{1})$ and $B(\mathcal{H}_{2})$ respectively represent the space of bounded linear operators on them. Let $\{e_{ij}\}_{i,j=1}^{n}$ and $\{f_{ij}\}_{i,j=1}^{m}$ respectively represents the complete set of matrix units for $B(\mathcal{H}_{1})$ and $B(\mathcal{H}_{2})$. Let $\Phi:B(\mathcal{H}_{1})\rightarrow B(\mathcal{H}_{2})$ be a linear map. Then the Choi matrix [5, 6, 1, 7, 8] of $\Phi$ is defined as
$C_{\Phi}=\sum_{{i,j}=1}^{n}e_{ij}\otimes\Phi(e_{ij})\in B(\mathcal{H}_{1})\otimes B(\mathcal{H}_{2}).$ (13)
The map $\Phi\mapsto C_{\Phi}$ is known as the CJKS isomorphism with the inverse given by

### v2

@@ 2.1.1 Partial Ordering and Möbius Inversion
The incidence algebra is a unital algebra with the identity element given by $\mathrm{I}(x,y)=\delta_{xy}$, where $\delta_{xy}$ denotes the Kronecker delta. With the identity being present in the incidence algebra, the inverse of $f\in R[S\times S]$ can be defined if there exist $f^{-1}\in R[S\times S]$ such $f*f^{-1}(x,y)=f^{-1}*f(x,y)=\mathrm{I}(x,y)$ for all $x,y\in S$ with $x\leq y$. Equivalently,
@@
The above equation is used to find $f^{-1}(x,y)$ inductively. Specifically, for all $x\in S$, we have $f^{-1}(x,x)=\frac{1}{f(x,x)}$. Then using this value one can compute $f^{-1}(x,z)$ for those elements $z\in S$ such that there is no $z_{0}\in S$ with $x<z_{0}<z$. Proceeding inductively in this way, one can compute $f^{-1}(x,y)$.
Definition (Zeta function and Möbius function [27]).
The zeta function is the element $\zeta\in R[S\times S]$ defined by
$\zeta(x,y)=\begin{cases}1,&\text{if}\quad x\leq y,\\ 0,&\text{otherwise.}\end{cases}$
The Möbius function, denoted by $\mu$, is the inverse of the zeta function in the incidence algebra.
It is clear that $\mu(x,x)=\frac{1}{\zeta(x,x)}=1$. Moreover, when $x<y$, from Eq. (3) we have
@@
Equation (4) provides a way to inductively determine the Möbius function.
Theorem B (Möbius inversion formula [27]).
Let $S$ be a partially ordered set with the ordering relation $\leq$, and let $R$ be a field. Suppose $f,h:S\to R$ are functions satisfying, for all $x\in S$,
@@
Then $h$ can be expressed in terms of $f$ as
@@
2.1.2 Steinberg’s Isomorphism and Representations of the Contracted Algebra of Finite Inverse Semigroups
An example of an inverse semigroup is the symmetric inverse semigroup. Denoted by $\mathcal{J}_{X}$, it consists of all partial bijections $\tau:Y\to Z$ on a non-empty set $X$, where $Y,Z\subseteq X$. The binary operation is the composition of partial transformations. It can be verified that $\tau:Y\to Z$ and its inverse $\tau^{-1}:Z\to Y$ satisfy all the defining properties of an abstract inverse semigroup as given in Definition A. The Preston-Vagner representation theorem states that any inverse semigroup $S$ can be thought of as an inverse subsemigroup of $\mathcal{J}_{S}$.
Theorem (Preston-Vagner representation [26, 12, 28]).
Let $S$ be an inverse semigroup, and let $\mathcal{J}_{S}$ denote the symmetric inverse semigroup of all partial bijection of $S$. Then $S$ is isomorphic to an inverse subsemigroup of $\mathcal{J}_{S}$.
For $\tau\in\mathcal{J}_{S}$, denote by $\mathrm{dom}(\tau)$ the domain of $\tau$ and by $\mathrm{ran}(\tau)$ its range. Denoting by $\mathrm{id}_{U}$ the identity map on a subset $U\subseteq S$, we have $\tau\tau^{-1}=\mathrm{id}_{\mathrm{ran}(\tau)}$ and $\tau^{-1}\tau=\mathrm{id}_{\mathrm{dom}(\tau)}$. Since $\mathrm{id}_{U}$ can be identified with the subset $U\subseteq S$ itself, it is customary to define, for an abstract inverse semigroup $S$ and $s\in S$,
$\mathrm{dom}(s):=s^{-1}s,\quad\mathrm{ran}(s):=ss^{-1},$ (7)
and to regard $s$ as an isomorphism $s:s^{-1}s\to ss^{-1}$. Recalling that $s^{-1}s$ and $ss^{-1}$ are idempotents, the notion of two idempotents being isomorphic is defined as follows.
Definition (Isomorphic idempotents [14]).
Let $S$ be an inverse semigroup. Two idempotents $e,e^{\prime}\in S$ are said to be isomorphic if there exists an element $s\in S$ such that
$e=s^{-1}s\quad\text{and}\quad e^{\prime}=ss^{-1}.$
@@ 2.1.2 Steinberg’s Isomorphism and Representations of the Contracted Algebra of Finite Inverse Semigroups
Definition C (Partial order [12, 28]).
Let $S$ be an inverse semigroup and let $s,t\in S$. We define $s\leq t$ if there exist an idempotent $e\in S$ such that $s=et$.
For two arbitrary elements $s,t\in S$ of an inverse semigroup $S$, the following statements are equivalent:
(i) $s\leq t,$
(ii) $\text{there exists and idempotent }e^{\prime}\text{ such that }s=te^{\prime},$
(iii) $s=ss^{-1}t,$
(iv) $s=ts^{-1}s.$
Definition D (Groupoid basis [14] ).
Let $S$ be a finite inverse semigroup. For every non-zero element $s\in S$, define
$\lfloor s\rfloor:=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t,s)t\in\mathbb{C}_{0}[S],$ (8)
where $\mu(t,s)$ is the Möbius function defined on the natural partial order on $S$, as given in Eq. (4).
It immediately follows that for the zero element $z\in S$, $\lfloor z\rfloor=0$. It is proved in Ref. [14] that $\{\lfloor s\rfloor\ :s\in S\setminus\{z\}\}$ forms a basis for the contracted algebra $\mathbb{C}_{0}[S]$. The basis $\{\lfloor s\rfloor\}_{s\in S\setminus\{z\}}$ is known as the groupoid basis. Using the Möbius inversion formula, the natural basis $\{s\}_{s\in S\setminus\{z\}}$ can be recovered as follows:
$s=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\lfloor t\rfloor.$ (9)
The basis elements in $\{\lfloor s\rfloor\}_{s\in S\setminus\{z\}}$ multiply as follows [14]:
@@
Let $S$ be a finite inverse semigroup. For $s,t\in S$, we say that $s$ and $t$ are $\mathcal{D}$-related, written $s\mathcal{D}t$, if there exists an element $x\in S$ such that $x^{-1}x=ss^{-1}$ and $xx^{-1}=tt^{-1}$.
@@
The relation defined above is called Green’s $\mathcal{D}$-relation. It can be shown that this relation is an equivalence relation on $S$ and its equivalence classes are referred to as the $\mathcal{D}$-classes. Let $\{D_{k}\}_{k=1}^{n}$ denote the $\mathcal{D}$-classes of $S$. Define
$\mathbb{C}D_{k}:=\mathrm{Span}\{\lfloor s\rfloor:s\in D_{k}\}.$
@@
It should be noted that there always exists an idempotent inside each $\mathcal{D}$-classes $D_{k}$. Indeed for any $t\in D_{k}$, one has $t^{-1}t\in D_{k}$. Moreover, if $e$ and $f$ are two idempotents in $D_{k}$, then they are isomorphic, i.e., there exists $x\in S$ such that $x^{-1}x=e$ and $xx^{-1}=f$.
Proposition E.
For each $\mathcal{D}$-class $D_{k}$, the vector space $\mathbb{C}D_{k}$ is a subalgebra of $\mathbb{C}_{0}[S]$.
@@
Proof: From the definition, $\mathbb{C}D_{k}$ is a vector space. Let $s,t\in D_{k}$ be two arbitrary elements. So, the multiplication of $\lfloor s\rfloor$ and $\lfloor t\rfloor$ is given by Eq. (10). It is sufficient to show that $st\in D_{k}$. Since $s,t\in D_{k}$, there exist $x\in S$, such that $x^{-1}x=ss^{-1}$ and $xx^{-1}=tt^{-1}$. Now,
$(st)(st)^{-1}$ $=stt^{-1}s^{-1}=ss^{-1}ss^{-1}=ss^{-1}=x^{-1}x,$
@@
where in the 2nd equality, we have used the condition in Eq. (10). It follows from the above equation that $(st)\mathcal{D}t$ and hence $st\in D_{k}$. This completes the proof. $\Box$
@@
A subset $G\subseteq S$ is called a subgroup of a semigroup $S$ if $G$ forms a group under the operation in $S$. A subgroup $G$ of $S$ is called a maximal subgroup if it is not properly contained in any other subgroup of $S$.
Proposition ([14]).
@@
Let $S$ be a finite inverse semigroup, and let $e\in S$ be an idempotent. Then $e$ is the identity element of a unique maximal subgroup $G_{e}\subseteq S$ given by
$G_{e}=\{s\in S:ss^{-1}=s^{-1}s=e\}.$
For an idempotent $e\in S$, the subgroup $G_{e}$ is called the maximal subgroup at $e$. The following proposition states that maximal subgroups at different idempotents belonging to the same $\mathcal{D}$-class are isomorphic.
Proposition ([13, 14]).
@@
Let $e$ and $f$ be two isomorphic idempotents of a finite inverse semigroup $S$, and let $G_{e}$ and $G_{f}$ denote their maximal subgroups. Then $G_{e}$ is isomorphic to $G_{f}$.
From Proposition E and the multiplication rule given in Eq. (10), it follows that
$\mathbb{C}_{0}[S]=\bigoplus_{k=1}^{n}\mathbb{C}D_{k}.$
@@
For each $D_{k}$, let us fix an idempotent $e_{k}$. Let $G_{e_{k}}$ denotes the maximal subgroup of $S$ at $e_{k}$. For each idempotent $e\in D_{k}$, let us fix an element $p_{e}\in S$ such that $\mathrm{dom}(p_{e})=e_{k}$ and $\mathrm{ran}(p_{e})=e$ with $p_{e_{k}}=e_{k}$. These choices will be fixed throughout the remainder of the paper. The following theorem states that $\mathbb{C}D_{k}$ is isomorphic to $M_{r_{k}}(\mathbb{C}G_{e_{k}})$, where $M_{r_{k}}(\mathbb{C}G_{e_{k}})$ is the algebra of $r_{k}\times r_{k}$ matrices over the group algebra $\mathbb{C}G_{e_{k}}$ with $r_{k}$ being the number of idempotents in $D_{k}$.
Theorem F ([14]).
Let $S$ be a finite inverse semigroup, and let $\{D_{k}\}_{k=1}^{N}$ denote its $\mathcal{D}$-classes. For each $k$, fix an idempotent $e_{k}\in D_{k}$, and let $G_{e_{k}}$ be the maximal subgroup of $S$ at $e_{k}$. Then
$\mathbb{C}_{0}[S]\cong\bigoplus_{k=1}^{N}M_{r_{k}}(\mathbb{C}G_{k}).$
Specifically, The algebra $\mathbb{C}D_{k}$ is isomorphic to the algebra $M_{r_{k}}(\mathbb{C}G_{k})$, where the isomorphism is given by
$\varphi(\lfloor s\rfloor)=p^{-1}_{\mathrm{ran}(s)}s~p_{\mathrm{dom}(s)}E_{\mathrm{ran}(s),\mathrm{dom}(s)},$ (11)
@@
$\varphi^{-1}(sE_{t,f})=\lfloor p_{t}sp^{-1}_{f}\rfloor,$ (12)
where $E_{a,b}$ is the $r_{k}\times r_{k}$ matrix indexed by the idempotents of $D_{k}$ whose $a,b$-th entry is $1$ and the rest of the entries are zero.
Proposition G.
Let $s\in D_{k}$. Then $p^{-1}_{\mathrm{ran}(s)}s~p_{\mathrm{dom}(s)}\in G_{e_{k}}$.
Proof: We begin by evaluating $\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)^{-1}\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)$.
@@
In a similar computation shows that
$\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)\left(p^{-1}_{\mathrm{ran}(s)}sp_{\mathrm{dom}(s)}\right)^{-1}=e_{k}.$
Thus $p^{-1}_{\mathrm{ran}(s)}s~p_{\mathrm{dom}(s)}$ is an element of the maximal subgroup $G_{e_{k}}$. This completes the proof. $\Box$
Proposition H.
For all $u\in G_{e_{k}}$ and idempotents $a,b\in D_{k}$, we have
$\mathrm{ran}(p_{a}\,u\,p^{-1}_{b})=a\quad\text{and}\quad\mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b.$
Proof: We compute:
$\mathrm{ran}(p_{a}\,u\,p^{-1}_{b})$ $=p_{a}\,u\,p^{-1}_{b}\,p_{b}\,u^{-1}\,p^{-1}_{a}$
@@
A similar computation shows that
$\mathrm{dom}(p_{a}\,u\,p^{-1}_{b})=b.$
This completes the proof. $\Box$
Proposition I.
Let $S$ be a finite inverse semigroup and let $s,t\in S$ be nonzero elements satisfying $ss^{-1}=tt^{-1}$. Then $s^{-1}t$ is an idempotent if and only if $s=t$.
Proof: When $s=t$, trivially $s^{-1}t$ is an idempotent. Conversely, suppose that $s^{-1}t$ is an idempotent. Then
$s^{-1}ts^{-1}t=s^{-1}t,$ (13)
$s^{-1}t=t^{-1}s,$ (14)
and
$(s^{-1}t)^{-1}(s^{-1}t)=(s^{-1}t)(s^{-1}t)^{-1}.$ (15)
Using Eq. (15) and the assumption $ss^{-1}=tt^{-1}$, we have
$s^{-1}s=t^{-1}t.$ (16)
Now we use Eq. (14) in Eq. (13) to obtain
$s^{-1}tt^{-1}s=s^{-1}t\iff s^{-1}ss^{-1}s=s^{-1}t\iff s^{-1}s=s^{-1}t\iff ts^{-1}s=ts^{-1}t.$
Now, using Eq. (16), we arrive at
$tt^{-1}t=ts^{-1}t,$ (17)
which is equivalent to
$t=ts^{-1}t.$ (18)
Since in a finite inverse semigroup every element has a unique inverse, we conclude from Eq. (18) that $s=t$. This completes the proof. $\Box$
Steinberg’s isomorphism provides a method for constructing a complete set of inequivalent irreducible representations of the contracted algebra $\mathbb{C}_{0}[S]$ from complete sets of inequivalent irreducible representations of the maximal subgroups of a finite inverse semigroup $S$.
Theorem J ([14]).
Let $S$ be a finite inverse semigroup, and let $\{D_{k}\}_{k=1}^{N}$ denote its $\mathcal{D}$-classes. For each $k$, fix an idempotent $e_{k}\in D_{k}$, and let $G_{e_{k}}$ be the maximal subgroup of $S$ at $e_{k}$. Let $\mathrm{Irr}(G_{e_{k}})$ be a complete set of inequivalent irreducible representations of the maximal subgroup $G_{e_{k}}$. For $\rho_{k}\in\mathrm{Irr}(G_{e_{k}})$, define a representation $\overline{\rho}_{k}$ of $M_{r_{k}}(\mathbb{C}G_{e_{k}})$ by
$\overline{\rho}_{k}(gE_{a,b}):=E_{a,b}\otimes\rho_{k}(g),$
extended linearly to the whole algebra. Now define $\overline{\rho}_{k}$ on the algebra $\bigoplus_{k=1}^{N}M_{r_{k}}(\mathbb{C}G_{e_{k}})$ by defining its action on the summands other than $k$ as zero. Then the set of all such $\overline{\rho}_{k}$ forms a complete set of inequivalent irreducible representations of the contracted semigroup algebra $\mathbb{C}_{0}[S]$. Specifically, the representation $\overline{\rho}_{k}$ is given on the groupoid basis by
$\overline{\rho}_{k}(\lfloor s\rfloor)=\begin{cases}E_{\mathrm{ran}(s),\mathrm{dom}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)})&\text{if}\quad s\in D_{k},\\ 0&\text{otherwise}.\end{cases}$ (19)
2.2 Choi-Jamiołkowski Isomorphism and Completely Positive Maps
Let $\mathcal{H}_{1}\!=\!\mathbb{C}^{n}$ and $\mathcal{H}_{2}=\mathbb{C}^{m}$ be two finite-dimensional Hilbert spaces, and let $\mathcal{B}(\mathcal{H}_{1})$ and $\mathcal{B}(\mathcal{H}_{2})$ respectively denote the algebras of bounded linear operators on them. Let $\{e_{ij}\}_{i,j=1}^{n}$ and $\{f_{ij}\}_{i,j=1}^{m}$ respectively denote the complete set of matrix units for $\mathcal{B}(\mathcal{H}_{1})$ and $\mathcal{B}(\mathcal{H}_{2})$. Given a linear map $\Phi:\mathcal{B}(\mathcal{H}_{1})\rightarrow\mathcal{B}(\mathcal{H}_{2})$, its Choi matrix [29, 30, 15, 31, 32] is defined by
$C_{\Phi}=\sum_{{i,j}=1}^{n}e_{ij}\otimes\Phi(e_{ij})\in\mathcal{B}(\mathcal{H}_{1})\otimes\mathcal{B}(\mathcal{H}_{2}).$ (20)
The map $\Phi\mapsto C_{\Phi}$ is known as the Choi-Jamiołkowski isomorphism. Its inverse is given by